OpenStudy (anonymous):

Find dy: (x)/√(7x-4). Please show steps!

OpenStudy (australopithecus):

multiply the top and the bottom by the reciprical (7x+4)^(1/2)

OpenStudy (australopithecus):

do they want the derivative?

OpenStudy (anonymous):

Yes, the answer is 7x-8/2(7x-4)^(3/2( i don't know how they got 7x-8!! :(

OpenStudy (australopithecus):

do you want to use quient rule or product rule?

OpenStudy (australopithecus):

I prefer product rule

OpenStudy (anonymous):

It doesn't really matter, since it's practically the same thing!

OpenStudy (australopithecus):

yeah but some teachers are nazis about use of quient rule for questions such as these so we have f(x) = (x)(7x-4)^(-1/2) g(x) = x g'(x) = 1 s(x) = (7x-4)^(-1/2) s'(x) = -1/(2(7x-4)^(3/2)) * (7) CHAIN RULE to find s'(x) d(x) = x^(-1/2) d'(x) = -1/(2x^(3/2)) l(x) = 7x-4 l'(x) = 7 s'(x) = d'(l(x)) * l'(x) so now we apply product rule g'(x)s(x) + g(x)s'(x) so (7x-4)^(-1/2) + x(-1/(2(7x-4)^(3/2)) * (7))

OpenStudy (australopithecus):

best to do derivatives the long way like I have when you first start out until you can do them in your head

OpenStudy (australopithecus):

just split things into smaller functions and apply the rules

OpenStudy (anonymous):

Yes, I got to this part: (7x-4)^(-1/2) + x(-1/(2(7x-4)^(3/2)) * (7)) but how did they get 7x-8 in the numerator?

OpenStudy (anonymous):

What I'm trying to ask is how did they combine those two?

OpenStudy (australopithecus):

you know about negative exponents right?